OpenStudy (anonymous):

Hi, this is my first question here. How do you solve: 1/(z-i) + (2+i)/(1+i) =√2? Given clue: Sub z=a+bi Thanks!

OpenStudy (anonymous):

It look like we need get all this into the form of f(z) I can start for you 1/z-i + 2+i/1+i =√2 1/z-i =√2 - [2-i/1-i] . . .z-i = . z =